# Using Half-Precision for GNN Training

### v1

@@ Abstract.
Recent trends in lower precision, e.g. half-precision floating point, training have shown improved system performance and reduced memory usage for Deep Learning while maintaining accuracy. However, current GNN systems cannot achieve such goals for GNN, as our analyses show that they massively underperform while showing abnormal accuracy when using half-precision. These systems suffer from value overflow issues due to lowered precision, under-utilization of hardware resources, and poor training performance. To mitigate this, we introduce HalfGNN, a half-precision based GNN system. HalfGNN proposes novel techniques: new vector operations for half-precision data types that improve data load and reduction performance, and discretized SpMM that overcomes the value overflow and natively provides workload balancing. Such techniques improve hardware utilization, reduce memory usage, and remove atomic writes. Evaluations show that HalfGNN achieves on average of 2.30×2.30$\times$2.30 × speedup in training time over DGL (float-based) for GAT, GCN, and GIN respectively while achieving similar accuracy, and saving 2.67$\times$ memory.
@@ 1. Introduction
Graph neural networks (GNN) have become ever more present in many deep learning applications  (xu2018powerful, ; courbariaux2015binaryconnect, ; han2020graph, ; krahmer2003graph, ; wang2021relational, ; wu2022graph, ; ying2018graph, ; zitnik2018modeling, ). This is because graphs can model different kinds of phenomena across various domains. It has been shown that GNNs offer better performances than other approaches (hamilton2017inductive, ; xu2018powerful, ) for graph datasets. GNN relies on two key sparse kernels: SpMM (Sparse matrix-matrix multiplication), and SDDMM (Sampled dense-dense matrix multiplication) and their many variants. These two kernels are very different from a GeMM (General Matrix Multiplication) kernel due to the unstructured sparsity pattern present in the graph dataset.
@@
GPU memory and computation capability have increased in recent years. However, it is still limited compared to the requirements of training large models. So, recent DL systems have explored lower precision floating-point data types such as half precision (16-bit floating point) instead of float (32-bit floating point) to utilize GPU resources efficiently. The half data type can theoretically provide faster computation and lower memory consumption than float. PyTorch (pytorch2019, ), a popular DL framework, has introduced mixed-precision training to integrate half data types. DGL (dgl2019, ), a popular GNN library, does provide support for half, which internally uses the Cusparse library for SpMM kernel, and in-house SDDMM. However, as we explain next, GNN remains to be explored using half precision.
@@
Training Runtime Efficiency. We plotted the SpMM runtime using Cusparse, which DGL uses internally, in Fig. 1(a) for different feature lengths on two widely used datasets in GNN literature. The dataset and machine configurations are discussed in  section 5. The results show that half-precision SpMM is significantly slower than single-precision float. Similarly, Fig. 1(b) shows that DGL SDDMM does not improve runtime when the half-precision is used. These results indicate that the performance characteristics of sparse kernels for half-precision are not yet fully explored.
@@
The performance characteristic is further complicated due to the availability of more hardware choices, such as tensor core, a hardware unit within GPU that performs faster computation on half-precision. To apply tensor cores to sparse kernels, we need to convert the sparse matrix to dense tiles; however, a GNN dataset may not be made up of dense tiles due to its unstructured sparsity of graph datasets. Unfortunately, TC-GNN (wang2023tc, ), which postulates tensor cores can provide better SpMM performance, actually causes huge performance degradation as shown by numerous prior works (Graphpy, ; GNNBench, ).
@@
Training Accuracy. Most importantly, Fig. 1(c) shows that DGL achieves significantly lower training accuracy when using half for Ogb-Product and Reddit datasets. It would be acceptable if there were small losses in training accuracy due to loss in precision. Unfortunately, the training loss becomes NaN(not a number) after a couple of iterations. The observation begs the question: whether the model or the underlying system design and integration is responsible for significantly lower accuracy?
@@
Contribution. These observations motivate us to explore and understand the GNN accuracy and performance characteristics when using half-precision, resulting in the proposed HalfGNN, the first ever half-precision based GNN System that achieves accuracy equivalent to float-based GNN and achieve significant runtime performance. We emphasize that the CUDA core is still the best for the computation as the tensor core is not suited due to the unstructured sparsity present in the graph datasets. Hence borrowing the existing idea from float-based sparse kernels becomes relevant. Unfortunately, adopting prior techniques fails to bring better performance because of several issues: (a) data-load does not achieve full memory coalescing in GPU due to the usage of half-precision as each GPU warp can only issue 64 bytes of data-load and not 128 bytes; (b) half-based arithmetic can only achieve sub-optimal throughput thereby lowering the performance; (c) current float-based workload balancing techniques use atomic instruction for conflicting writes, which is very costly for half-precision based operation.
To this end, we propose the following contribution:
@@
$\bullet$ Adopting Half2 Data-Type for a Strong Baseline for Half-Precision Sparse Kernels. We adopt half2 data-type, a native feature of GPUs that treats two half-precision numbers as a single entity for memory operation (data-load and data-store instructions) and computations, to achieve full coalesced memory access and better arithmetic throughput. Adopting is a challenge that we overcome using Edge Feature Mirroring, Feature Padding for Odd Vertex Features, and introducing sub-warps to impart better thread utilization. This establishes an ideal baseline for half-precision kernels which we analyze and propose following techniques as they still fail to provide the envisioned performance and accuracy.
@@
$\bullet$ More Vector Types for Improving SDDMM Performance. We introduce grouping of $4$ and $8$ half values, called (half4 and half8) respectively to exploit the native float2 and float4 support of GPUs to increase data-load performance by being able to issue more feature load instructions before the implicit memory barrier is encountered in the inter-thread communication for SDDMM reduction.
@@
$\bullet$ Overflow and Discretized Reduction For SpMM. We analyze the root cause of NaN training error in GNNs. A vertex with a large neighborhood size, common in graphs, can produce INF (infinity) due to value overflow in SpMM reduction as half-precision has a limited numerical range. This INF then leads to NaN generation due to follow-up operations such as softmax involving arithmetic on two INF. However, GNN models indeed offer built-in mechanisms to protect against overflow such as degree-norm scaling in GCN which scales the output of SpMM back to half-precision range. But, merely the presence of such a mechanism is insufficient unless explored systematically. This work studies many such mechanisms across different GNN models and embeds them into system design through discretized reduction where only a finite batch of dot products are reduced and scaled (normalized) before handling the next batch of dot products, thus protecting the output against overflow.
@@
The discretization is naturally suited to provide workload balancing in SpMM as one is free to choose the batch size, thereby allocating an equal amount of computation cores to each discretization unit leading to workload balancing. We demonstrate an edge-parallel and a vertex-parallel SpMM to show the generality of the proposed discretization for workload balancing and propose non-atomic write through usage of staging buffer and a follow-up kernel.
@@
$\bullet$ Fear of Overflow and Frequent Data-Conversion. Further study shows that a naive mix-precision system performs excessive data-type conversion of tensors (float to half and vice-versa) or invokes many float-based kernels thereby not providing much benefits in moving to mix-precision training. E.g., the PyTorch exponential operation works in float due to fear of overflow. This operation is used in the GAT attention mechanism, introducing half to float on its input tensor and float to half on the output tensor. However, these conversions are not needed due to the built-in properties of the GAT attention mechanism that guarantee no overflow. This paper studies many such operations to truly incorporate half-precision.
@@
Evaluations show that HalfGNN achieves state-of-the-art accuracy performance, along with massive improvement in runtime performance for half-precision. Briefly, we reduce the runtime on average of 2.30×2.30$\times$2.30 × and required memory by 2.67×2.67$\times$2.67 × on average over DGL for GAT, GCN, and GIN respectively. Furthermore, HalfGNN achieved 7.12×7.12$\times$7.12 ×, and 22.89×22.89$\times$22.89 × speedup over Cusparse half SpMM, and DGL half SDDMM.
@@
We applied some of the performance optimizations of HalfGNN to Huang et al.  (huang2021understanding, ) and achieved 1.79×1.79$\times$1.79 × speedup for SpMM, validating the generality of the performance optimizations.
@@
The remainder of the paper is organized as follows. We present the necessary background in section 2 and analyze the current GNN systems for half-precision, discuss related works, and present an overview in section 3. Detailed design is presented in section 4, evaluations in section 5, and conclusion in section 6.
@@ 2. Background
GPU. A program written for GPU is called a kernel. Computation units in GPUs are known as Symmetric Multiprocessors(SM). SM provides a programmable hardware cache, called shared memory, which the kernel can explicitly use to write and read. Programmers create and configure Cooperative Thread Arrays (CTA) or thread-blocks that are mapped to SM by the GPU driver. In a CTA, threads are usually grouped into 32 threads, called a warp, that executes computation in SIMT (single instructions multiple threads). Kernels are generally designed to access memory in a contiguous manner for better data-load performance.
@@
Graph Storage Format and Tensor Types Fig. 2 shows a sample graph G = {V, E} where V is the vertex set and E is the edge set. We use graph theory and linear algebra terminologies (e.g., row, columns, and non-zero elements (NNZ)) to describe the graph and features of vertices and edges. The two most commonly used graph formats are Compressed Sparse Row (CSR) and Coordinate Objects (COO). COO uses two arrays of rowID and ColID where each $(rowID,colID)$ pair denotes an edge or non-zero element (NZE). The CSR also uses two arrays for its format: offset array points to the start of the neighborhood of each vertex or row. The degree defines the neighbor size of each vertex.
@@
We also classify tensors involved in kernels, as vertex-level or edge-level tensors. For a graph having $|V|$ vertices and $|E|$ edges, each vertex having $|F|$ features would have a feature matrix of shape $|V|\times|F|$ (vertex-level tensor). For edge features, we get $|E|\times 1$ tensor (edge-level tensor).
Half Precision and Vector Data Types. Half-precision floating point numbers imply 16-bit floating point numbers and hence support a lower range of absolute values: $2^{-14}$ to $(2-2^{-10}).2^{15}$ i.e. $6.1\times 10^{-5}$ to $6.5\times 10^{4}$, which is very small compared to the range of single-precision float data-type. Hence, any value greater than the upper bound overflows to INF, while smaller values underflow to zero. Modern GPUs have provided support for half-precision arithmetic.
Vector types allows grouping of various lengths of basic type for data-load and computation. For example, two half numbers could be packed in one half2 vector data type, whose size is 32 bits. GPUs support native arithmetic on the half2 data type. On the other hand, float2 and float4 pack two and four float data. However, these two vector types support native data-load, but not arithmetic.
2.1. Deep Learning Traning and Sparse Kernels
@@
Deep learning training process has two parts: forward and backward computation. The forward calls many kernels from the input layer to the output layer deriving prediction values. They are compared with the ground truth to calculate the loss. Backward uses the loss and invokes kernels from the output layer to the input layer to derive gradients and update the model parameters. The backward computation accesses state-tensor– results that are produced in the forward computation.
@@
For sparse matrix dense matrix multiplication (SpMM) ($Y\leftarrow A_{w}X$), its backward computation needs to include both the SpMM ($\delta X\leftarrow A_{w}^{T}\delta Y$) and sampled dense dense matrix multiplication (SDDMM) ($\delta W_{e}\leftarrow A\odot(\delta YX^{T})$), where $\delta$ represents gradients of the tensors. The paper deals with these 2 classes of sparse kernels.
SpMM is categorized into 2 types. SpMMve is the usual SpMM, shown in Fig. 2d, where the edges contains features, and hence $A_{w}$ implies a graph topology ($A$) plus an edge-level tensor ($W_{e}$). It is then multiplied to a dense tensor X of size $|V|\times|F|$. A special case is SpMMv where all edge features are treated as 1.0, hence we do not need to store an explicit $W_{e}$. In SDDMM, shown in Fig. 2e, X and Y (or $\delta Y$) are two vertex-level tensors of size $|V|\times|F|$ while $A$ is meant to produce $W_{e}$ only (features/gradients on edges).
@@
SpMMv is the only sparse kernel that is used in GCN (kipf2016semi, ) and GIN (xu2018powerful, ). Attention-based GNNs, such as GAT (gat18iclr, ) would require both SpMMve and SDDMM.
@@
Sparse kernels broadly have 3 key steps: a) Data Load brings graph and features from the GPU global memory to the GPU registers of the computation unit. b) Actual Computation performs parallel computations to compute dot product of features; c) Reduction and Data-Store first reduces the computed dot products followed by writing to the output (GPU) memory. SpMM and SDDMM vary in how they perform the reduction phase. SpMM reduction happens across neighborhood dimensions, needing inter-thread communication only when the neighborhood of a vertex is allocated more than one warp to process, as in workload-balanced SpMM. SDDMM performs reduction across feature dimensions, so reduction mostly needs inter-thread communication.
@@
Vertex-parallel (bell2009implementing, ) allocates equal computation to each vertex to process its neighborhood. Works such as (gale2020sparse, ; featgraphsc20, ; gespmmsc2020, ; fu2022tlpgnn, ; wu2021seastar, ) uses vertex-parallel variants. Workload imbalance occurs when certain vertices have large neighborhoods compared to others. Edge-parallel systems (yang2018design, ; gnnone, ) divide edges equally among computation units leading to a fully workload-balanced design. They introduce the need for inter-thread communication in SpMM for reduction and atomic write for data-store. Both types of parallel systems dedicate feature-parallel methods (yang2018design, ) when loading the vertex-level features leading to coalesced memory access for the same.

### v2

# Optimization of GNN Training Through Half-precision

@@ Abstract.
[…]
Graph Neural Networks, GPU, Performance
journalyear: 2025copyright: ccconference: The 34th International Symposium on High-Performance Parallel and Distributed Computing; July 20–23, 2025; Notre Dame, IN, USAbooktitle: The 34th International Symposium on High-Performance Parallel and Distributed Computing (HPDC ’25), July 20–23, 2025, Notre Dame, IN, USAdoi: 10.1145/3731545.3731575isbn: 979-8-4007-1869-4/2025/07ccs: General and reference Performanceccs: Computing methodologies Neural networks
@@ 1. Introduction
Graph neural networks (GNN) have become ever more present in many deep learning applications  (xu2018powerful, ; courbariaux2015binaryconnect, ; han2020graph, ; krahmer2003graph, ; wang2021relational, ; wu2022graph, ; ying2018graph, ; zitnik2018modeling, ; fan2019graph, ; hamilton2018embedding, ; zhang2018gaan, ). This is because graphs can model different kinds of phenomena across various domains. It has been shown that GNNs offer better performances than other approaches (hamilton2017inductive, ; xu2018powerful, ) for graph datasets. GNN relies on two key sparse kernels: SpMM (Sparse matrix-matrix multiplication), and SDDMM (Sampled dense-dense matrix multiplication) and their many variants. These two kernels are very different from a GeMM (General Matrix Multiplication) kernel due to the unstructured sparsity pattern present in the graph dataset.
@@
GPU memory and computation capability have increased in recent years. However, it is still limited compared to the requirements of training large models. So, recent DL systems have explored lower precision data types such as half precision (16-bit floating point) instead of float (32-bit floating point) to utilize GPU resources efficiently. The half data type can theoretically provide faster computation and lower memory consumption than float. PyTorch (pytorch2019, ), a popular DL framework, has introduced mixed-precision training to integrate half data types. DGL (dgl2019, ), a popular GNN library, does provide support for half, which internally uses the Cusparse (cusparse, ) library for the SpMM kernel, and in-house SDDMM. However, as we explain next, GNN remains to be explored using half precision.
@@
Training Runtime Efficiency. We plotted the SpMM runtime using Cusparse, which DGL uses internally, in Fig. 1(a) for different feature lengths on two widely used datasets in GNN literature. The dataset and machine configurations are discussed in  section 6. The results show that half-precision SpMM is significantly slower than float SpMM. Similarly, Fig. 1(b) shows that DGL SDDMM does not improve runtime when the half-precision is used. These results indicate that the performance characteristics of sparse kernels for half-precision have not yet been fully explored.
@@
The performance characteristic is further complicated due to more hardware choices, such as the tensor core– a hardware unit within the GPU that performs faster computation on half-precision. To apply tensor cores to sparse kernels, we need to convert the sparse matrix to dense tiles; however, a GNN dataset may not be made up of dense tiles due to its unstructured sparsity of graph datasets. Unfortunately, TC-GNN (wang2023tc, ), which postulates that tensor cores can provide better SpMM performance, causes huge performance degradation, as shown by numerous prior works (Graphpy, ; GNNBench, ).
@@
Training Accuracy. Most importantly, Fig. 1(c) shows that DGL achieves significantly lower training accuracy when using half-precision for Ogb-Product and Reddit datasets. It would be acceptable if there were small losses in training accuracy due to a loss in precision. Unfortunately, the training loss becomes NaN(not a number) after a couple of iterations, questioning whether the model or the underlying system design is responsible for significantly lower accuracy.
@@
Contribution. These observations motivate us to explore and understand the GNN accuracy and runtime characteristics when using half-precision, resulting in the proposed HalfGNN, the first ever half-precision based GNN System that achieves accuracy equivalent to float-based GNN and achieves significant runtime performance. We emphasize that the CUDA core is still the best for computation, as the tensor core is unsuited due to the unstructured sparsity in graph datasets. Hence, borrowing the existing idea from float-based sparse kernels becomes relevant.
Unfortunately, adopting prior float-based techniques fails to bring better performance because of several issues: (a) data-load does not achieve full memory coalescing in GPU due to the usage of half-precision as each GPU warp can only issue 64 bytes of data-load and not 128 bytes; (b) half-based arithmetic can only achieve sub-optimal throughput, thereby lowering the performance; (c) current float-based workload balancing techniques handle conflicting writes using atomic instructions, which is very costly for half-precision based operation.
To this end, we propose the following contributions:
@@
$\bullet$ Adopting Half2 Data-Type for a Strong Baseline for Half-Precision Sparse Kernels. We adopt half2 data-type. This native GPU vector data type combines two half-precision numbers for memory operations (data-load and data-store instructions) and computations to achieve full coalesced memory access and better arithmetic throughput. Its adoption faces many challenges that we overcome using Edge Feature Mirroring, Feature Padding for Odd Vertex Features, and sub-warps to impart better thread utilization. This establishes an ideal baseline for half-precision kernels, which we analyze, and then propose the following techniques as the baseline fails to provide the envisioned performance and accuracy.
@@
$\bullet$ More Vector Types for Improving SDDMM Performance. We group $4$ and $8$ half values, called (half4 and half8) vector data types respectively. The proposed vector data types increase data-load performance by being able to issue more feature load instructions before the implicit memory barrier is encountered in the inter-thread communication for SDDMM reduction. On the other hand, the proposed data types minimize the number of inter-thread communications required in SDDMM.
@@
$\bullet$ Value Overflow and Discretized Reduction For SpMM. We analyze the root cause of the NaN training error in GNNs. A vertex with a large neighborhood size, common in graphs, can produce INF (infinity) due to value overflow in SpMM reduction as half-precision has a limited numerical range. This INF then leads to NaN generation due to follow-up operations such as softmax involving arithmetic on two INF. However, GNN models indeed offer built-in mechanisms to protect against overflow, such as degree-norm scaling in GCN (kipf2016semi, ), which scales the output of SpMM back to half-precision range. However, the presence of such a mechanism is insufficient unless explored systematically. This work studies many such mechanisms across different GNN models and embeds them into system design through discretized reduction where only a finite batch of dot products are reduced and scaled (normalized) before handling the next batch of dot products, thus protecting the output against overflow.
@@
The discretization is naturally suited to provide workload balancing in SpMM as one is free to choose the batch size, thereby allocating an equal number of computation cores to each discretization unit, leading to workload balancing. We demonstrate an edge-parallel and a vertex-parallel SpMM to show the generality of the proposed discretization for workload balancing and propose non-atomic write through usage of staging buffer and a follow-up kernel.
@@
$\bullet$ Fear of Overflow and Frequent Data-Conversion. Our Further analysis shows that a naive mixed-precision system performs excessive data-type conversion of tensors (float to half and vice-versa) or invokes many float-based kernels, thereby not providing much benefit in moving to mixed-precision training. For example, the PyTorch exponential operation works in float due to the fear of overflow. This operation is used in the GAT attention mechanism, introducing half to float on its input tensor and float to half on the output tensor. However, these conversions are redundant due to the built-in properties of the GAT attention mechanism that guarantee no overflow. This paper studies many such operations to truly incorporate half-precision.
@@
Evaluations show that HalfGNN achieves state-of-the-art accuracy along with huge improvement in runtime performance for half-precision. Briefly, we reduce the runtime on average by $2.30\times$ and required memory by $2.67\times$ on average over DGL for GAT, GCN, and GIN, respectively. Furthermore, HalfGNN achieved $7.12\times$, and $22.89\times$ speedup over Cusparse half SpMM, and DGL half SDDMM.
@@
We applied some of the performance optimizations of HalfGNN to Huang et al.  (huang2021understanding, ) and achieved $1.79\times$ speedup for SpMM, validating the generality of the performance optimizations.
@@
The remainder of the paper is organized as follows. We present the necessary background in section 2 and analyze the current GNN systems for half-precision, discuss related works, and present an overview in section 3. Detailed design is in section 4 and section 5, evaluations in section 6, and conclude in section 7.
@@ 2. Background
A program written for a GPU is called a kernel. Computation units in GPUs are known as streaming multiprocessors (SM). SM provides a programmable hardware cache, called shared memory, which the kernel can explicitly use to write and read. Programmers create and configure Cooperative Thread Arrays (CTA) or thread-blocks that are mapped to SM by the GPU driver. In a CTA, threads are usually grouped into 32 threads, called a warp, which executes computation in SIMT (single instruction multiple threads). Kernels are generally designed to access memory in a contiguous manner for better data-load performance.
GPU has become a very important accelerator for training diverse deep-learning models. Given this background on GPU, we next describe the graph storage and GNN training steps, explaining the computation steps of major sparse kernels.
2.1. Graph Neural Network and Sparse Kernels
2.1.1. Graph Storage Format
@@
Fig. 2 shows a sample graph G = {V, E} where V is the vertex set and E is the edge set. We use graph theory and linear algebra terminologies (e.g., row, columns, and non-zero elements (NNZ)) to describe the graph and features of vertices and edges. COO (Coordinate Objects) storage format uses two arrays of rowID and ColID, where each $(rowID,colID)$ pair denotes an edge or non-zero element (NZE). The CSR (Compressed Sparse Row) storage format also uses two arrays: offset array points to the start of the neighborhood of each vertex or row. The degree defines the neighborhood size of each vertex.
@@
We classify tensors involved in kernels as vertex-level or edge-level tensors. For a graph having $|V|$ vertices and $|E|$ edges, each vertex having $|F|$ features would have a feature matrix of shape $|V|\times|F|$ (vertex-level tensor). We get $|E|\times 1$ tensor (edge-level tensor) for edge features.
@@
GNN training, like deep learning, has two parts: forward computation calls many kernels from the input layer to the output layer, deriving prediction values. They are compared with the ground truth to calculate the loss. Backward computation uses the loss and invokes kernels from the output layer to the input layer to derive gradients and update the model parameters. The backward computation accesses state-tensor– results that are produced in the forward computation.
2.1.2. Sparse Kernels
The paper deals with these 2 classes of sparse kernels: sparse matrix dense matrix multiplication (SpMM) and sampled dense dense matrix multiplication (SDDMM). SpMM ($Y\leftarrow A_{w}X$) is shown in Fig. 2d, where the edges contains features, and hence $A_{w}$ implies a graph topology ($A$) plus an edge-level tensor ($W_{e}$). It is then multiplied to a vertex-level tensor X of size $|V|\times|F|$, producing another vertex-level tensor $Y$ of size $|V|\times|F|$.
@@ 2.1.2. Sparse Kernels
Its backward computation needs to include both the SpMM ($\delta X\leftarrow A_{w}^{T}\delta Y$) and SDDMM ($\delta W_{e}\leftarrow A\odot(\delta YX^{T})$), where $\delta$ represents gradients of the tensors. SDDMM is shown in Fig. 2e produces $\delta W_{e}$ that represents gradients on edges. One can note that X, Y, and $W_{e}$ are state-tensors.
@@
SpMM is categorized into 2 types. SpMMve is the usual SpMM as described above. A special case is SpMMv where all edge features are treated as 1.0, hence we do not need to store an explicit $W_{e}$. SpMMv is the only sparse kernel that is used in GCN (kipf2016semi, ) and GIN (xu2018powerful, ). Attention-based GNNs, such as GAT (gat18iclr, ) would require both SpMMve and SDDMM.
@@
Sparse kernels broadly have 3 key steps. a) Data Load brings graph and features from the GPU memory to the registers of the computation unit. b) perform parallel dot products of features: between features of row and column ID for each NZE in case of SDDMM, and between features of NZE (i.e., edge weight) and features of column ID in SpMM. c) Reduction and Data-Store first reduces the computed dot products, followed by writing to the output (GPU) memory: SDDMM performs reduction across feature dimensions, so reduction mostly needs inter-thread communication and are stored for each edge; and SpMM reduction happens across neighborhood dimensions, needing inter-thread communication only when the vertex neighbors are allocated more than one warp to process, as in workload-balanced SpMM and are stored for each row ID.
2.1.3. Parallelism.
@@
Vertex-parallel (bell2009implementing, ) sparse kernels allocate equal computation to process the neighborhood of each vertex. For example, one warp is needed to process the neighborhood of each vertex. Works such as (gale2020sparse, ; featgraphsc20, ; gespmmsc2020, ; fu2022tlpgnn, ; wu2021seastar, ) uses vertex-parallel variants. Workload imbalance occurs when certain vertices have a larger neighborhood than others. Edge-parallel systems (yang2018design, ; gnnone, ) divide edges equally among computation, leading to a fully workload-balanced design. However, they introduce the need for inter-thread communication in SpMM for reduction and atomic write for the data-store phase.
Both systems use feature-parallel (yang2018design, ) method when loading the vertex-level features: 32 threads of the warp load 32 different features at once, but 2 features are shown in Fig. 2 for illustration of matrix $X$. This leads to coalesced memory access from $X$.
2.2. Half-Precision and Vector Data-Type
Half-precision floating point numbers imply 16-bit floating point numbers and hence support a lower range of absolute values: $2^{-14}$ to $(2-2^{-10}).2^{15}$ i.e. $6.1\times 10^{-5}$ to $6.5\times 10^{4}$, which is very small compared to the range of single-precision float data-type. Hence, any value greater than the upper bound overflows to INF, while smaller values underflow to zero. Modern GPUs have provided support for half-precision arithmetic.
Figure 3. Arithmetic operations on Half
If one relies on native operators, such as ‘+’ or ‘*’, the data is implicitly converted to float, where the operation is performed in float (Fig. 3a). One can use CUDA intrinsic APIs for half computation that avoids this data-type conversion but achieves the same throughput as the float (Fig. 3b).
Vector data types allow grouping of various lengths of basic types for the data-load and the computation phases. For example, two half numbers could be packed in one half2 vector data type, whose size is 32 bits. GPUs support native data-load and arithmetic on the half2 data type (Fig. 3c), though DGL does not use it in their kernels.
Kindly note that the GPU supports float2 (64 bits) and float4 (128 bits) that pack two and four float data, respectively. They support native data-load, but not arithmetic. For half-precision, beyond half2 (32 bits), no other half-based vector data types are supported. This paper explores additional vector types for half-precision.